OpenStudy (anonymous):

What is the diameter of the circle if the radius is 5.7

OpenStudy (ajprincess):

diameter=2*radius

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large 2r = d }\) Like aj said

OpenStudy (anonymous):

Im sorry I mean the circumfrence

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large 2 \times \pi \times radius }\)

OpenStudy (anonymous):

what?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 6.28 \times 5.7 }\)

Parth (parthkohli):

Now a calculator ;p

OpenStudy (anonymous):

why 6.28

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2 \times 3.14 \times 5.7 }\) \( \color{Black}{\Rightarrow 6.28 \times 5.7 }\)

OpenStudy (anonymous):

okay , I get it now thank you

OpenStudy (anonymous):

thanks